OpenStudy (anonymous):

#5 125 ^4/3 = 625 A. log ^4/3 625=125 B. 3log(base 4)625=125 C. log (base125)625=4/3 D log (base625)125 =3/4

OpenStudy (anonymous):

used to hate these until I learned logs. answer = c

OpenStudy (anonymous):

\[\log _{b}x =y \]

OpenStudy (anonymous):

or \[b ^{y}=x\]

OpenStudy (anonymous):

so: 125^4/3=x corresponds to b^y=x and log(base125)625=4/3

OpenStudy (anonymous):

hope that helps!